# Training-Free Message Passing for Learning on Hypergraphs

### v4

@@ 3.3 Training-Free Message Passing
Learnable parameters removal. Based on the insights from Subsection 3.2, to develop a training-free MP-Module, we first remove learnable parameters from the feature aggregation functions of the selected HNNs. Specifically, we replace the learnable matrices in these functions with identity matrices. This removal allows us to reformulate the feature aggregation functions as follows:
@@
Linearisation. Inspired by the insights from Subsection 3.2, we further remove the non-linearity from the functions represented by Eqs. (4$\sim$4d), thereby transforming an MP-Module with $L$ feature aggregation functions into a more efficient single propagation step. For the sake of brevity, we present the simplification results as Proposition 3.2 and provide its mathematical proof in Appendix D.
@@
Let $\mathbf{X}_{\mathcal{V}}$ be input node features, $\hat{\mathbf{X}}_{\mathcal{V}}$ be the output of the MP-Module of an UniGCNII/AllDeepSets/ED-HNN/Deep-HGNN with $L$ MP layers, and $\alpha\!\in\![\!0,1\!)$. Assume that learnable parameters and the non-linearity are removed from the module. For any hypergraph =$\mathcal{H}\!=\!\!\{\!\mathcal{V},\!\mathcal{E},\!\mathbf{H}\!\}$, there exists a clique expansion $\mathcal{G}\!=\!\{\!\mathcal{V},\mathbf{W}\!\}$ to make $\hat{\mathbf{X}}_{\mathcal{V}}\!=\!\Big{(}(1\!-\!\alpha)^{L}\!\mathbf{W}^{L}\!+\!\alpha\!\sum_{l=0}^{L-1}(1\!-\!\alpha)^{l}\!\mathbf{W}^{l}\Big{)}\mathbf{X}_{\mathcal{V}}$.
Based on Proposition 3.2, we design the formulation of our TF-MP-Module as the following equation:
@@
where $\delta(\cdot)$ is a function that returns 1 if $e_{k}$ connects $v_{i}$ and $v_{j}$ and returns 0 otherwise. The main intuition behind the design of this edge weight is that, under certain conditions, the value of $\mathbf{W}_{H_{ij}}$ is positively correlated with the probability of nodes $v_{i}$ and $v_{j}$ having the same label. Consequently, this edge weight design can make the weighted clique expansion fit the homophily assumption, which is that connected nodes tend to be similar to each other (McPherson et al., 2001). We theoretically elaborate on this intuition in Appendix F. Furthermore, inspired by previous works in the graph domain (Wu et al., 2020), we generate $\mathbf{S}$ based on a symmetrically normalised $\mathbf{W}_{H}$ with the self-loop. Specifically, we generate the $\mathbf{S}$ by:
@@ 4 Theoretical Analysis
Let $H_{0}^{v^{L}_{i}}$ be the entropy of information used by an HNN with $L$ feature aggregation functions in generating the features of a node $v_{i}$, $H_{1}^{v^{L}_{i}}$ be the entropy of information used by an TF-HNN with an $L$-layer TF-MP-Module for the same purpose, and $H_{2}^{v^{L}_{i}}$ denote the entropy of information within node $v_{i}$ and its $L$-hop neighbours on the hypergraph. Then, $H_{0}^{v^{L}_{i}}=H_{1}^{v^{L}_{i}}=H_{2}^{v^{L}_{i}}$.
This proposition theoretically demonstrates that TF-HNN is as effective as existing HNNs in utilising information from pre-defined hypergraphs to generate node features. Specifically, both TF-HNN and existing HNNs achieve this by aggregating neighbourhood information.
Robustness against the oversmoothing issue. According to Proposition 3.3, TF-HNN can use global interactions within the given hypergraph to generate node features by deepening the TF-MP-Module. However, as shown in a recent study (Chen et al., 2022), the message-passing-based models on hypergraphs may suffer from the oversmoothing issue. This issue refers to the tendency of a model to produce indistinguishable features for nodes in different classes as the model depth increases, which consequently degrades the performance of the framework in downstream tasks. Consequently, to further support the effectiveness of TF-HNN, we theoretically analyse its robustness against the oversmoothing issue. Inspired by previous works in the graph domain (Eswaran et al., 2020; Yang et al., 2021; Zhang et al., 2020), we show Proposition 4.2, with the proof provided in Appendix H.
@@
Based on Proposition 4.2, a deep TF-MP-Module tends to assign similar features to connected nodes on a hypergraph while maintaining the unique input information of each node. Hence, even with an extremely large number of layers, the TF-MP-Module can preserve the unique input information of each node, making it robust against the oversmoothing issue. This property further ensures the effectiveness of our TF-HNN. Section 6 empirically supports the analysis presented in this section.
@@ 5 Related Work
Recent efforts have leveraged hypergraphs to improve node feature learning for downstream tasks (Bick et al., 2023). The predominant models in this area are hypergraph neural networks (HNNs) with a message passing module that enables information exchange between connected nodes (Antelmi et al., 2023). However, as noted for graph counterparts (Wu et al., 2019; Frasca et al., 2020) for their graph counterparts, HNNs suffer from low learning efficiency. Two recent works attempt to mitigate this limitation (Feng et al., 2024; Tang et al., 2024). In Feng et al. (2024), an HNN named HGNN (Feng et al., 2019) is distilled into an MLP during training; while in (Tang et al., 2024), hypergraph structural information is integrated into an MLP via a loss function computed by the sum of the maximum node feature distance in each hyperedge. These approaches primarily focus on reducing model inference complexity via a dedicated training process. However, these designs still lead to hypergraph machine learning models with computationally intensive training procedures. For instance, the distillation process in (Feng et al., 2024) necessitates training a HNN with a computational complexity of $\mathcal{O}(M+TLm^{\prime}d+TLnd^{2})$, where with $M$ is the training computational complexity of the task-specific module, $n$ is the node count, $m^{\prime}$ is the number of edges in the clique expansion, $L$ is the number of layers, and $d$ is the feature dimension. Given that many existing models are applied to semi-supervised hypergraph node classification settings, which involve the simultaneous use of training and test data (Huang & Yang, 2021; Chien et al., 2022; Wang et al., 2023a; b; Joachims et al., 1999; Chapelle et al., 2009), reducing the training complexity of hypergraph-based models remains a critical yet underexplored challenge. We address this challenge with a novel model named TF-HNN, which is the first to decouple the processing of hypergraph structural information from model learning. This decoupling significantly improves training efficiency.
@@ 6.1 Experimental Setup
Dataset and baseline. We conduct experiments on seven real-world hypergraphs: Cora-CA, DBLP-CA, Citeseer, Congress, House, Senate, and Trivago. We use nine HNNs as baselines. Five of these baselines utilize direct feature aggregation, including HGNN (Feng et al., 2019), HCHA (Bai et al., 2021), UniGCNII (Huang & Yang, 2021), PhenomNN (Wang et al., 2023b), and Deep-HGNN (Chen et al., 2022). The remaining four baselines employ indirect feature aggregation, consisting of HNHN (Dong et al., 2020), AllDeepSets (Chien et al., 2022), AllSetTransformer (Chien et al., 2022), and ED-HNN (Wang et al., 2023a). The dataset statistics are summarized in Table 2.
@@
Implementation. For node classification, we follow previous works (Wang et al., 2023a; Duta et al., 2023) to use a 50%/25%/25% train/validation/test data split and adapt the baseline classification accuracy from them [^3] Since there are no reported results for Deep-HGNN under the chosen data split, we use our own results.. Additionally, similar to these works, we implement the classifier based on MLP for our TF-HNN and report the results from ten runs. For hyperlink prediction, existing works (Chen & Liu, 2023) primarily focus on the design of the prediction head, with no reported results for applying our baseline HNNs to this task; so we report all results from five runs conducted by ourselves. We employ the deep set function implemented by (Chien et al., 2022) as the prediction head for our method and baseline methods due to its simplicity. We use a 50%/25%/25% train/validation/test data split and ensure each split contains five times as many fake hyperedges as real hyperedges. TF-HNN and HNNs only use real hyperedges in training and validation sets to do message passing. All experiments were on an RTX 3090 with PyTorch, and our code with running commands is at https://anonymous.4open.science/r/TF-HNN. More details are in Appendix L.
@@ 6.2 Comparison with Baselines
Node classification. We summarize the classification accuracy and relative training time of the TF-HNN and HNNs in Table 3 and Figure 2, respectively. Table 3 demonstrates that TF-HNN not only leads to the best results across four datasets (Cora-CA, Congress, House, and Senate) but also results in the highest average mean accuracy overall. These findings confirm the effectiveness of TF-HNN in generating powerful node features for node classification. As discussed in Section 4, we attribute this effectiveness to the ability of TF-HNN to match the power of existing HNNs in utilising neighbourhood information to generate node features, while requiring the optimization of fewer parameters. This makes the TF-HNN more efficient in utilising training data and reduces the risk of overfitting, thereby outperforming more complex HNN counterparts. On the other hand, as shown in Figure 2, the TF-HNN consistently needs less training time to achieve superior performance compared to the HNNs. This highlights the high training efficiency of TF-HNN in downstream tasks.
Hyperlink prediction. Figure 3 shows the results for hyperlink prediction, which exhibit a pattern similar to that observed in the hypergraph node classification task: the TF-HNN outperforms the HNNs while requiring less training time. An interesting observation is that the training efficiency improvement from TF-HNN is smaller in the hyperlink prediction task compared to the node classification task. The intuition is that the task-specific module used for hyperlink prediction is more computationally expensive than the one used for hypergraph node classification. Specifically, the hyperlink predictor needs to generate a prediction for each hyperedge by aggregating the features of the nodes connected by the hyperedge, whereas the node classifier only needs to perform forward propagation for each individual node. Appendix M shows that the efficiency improvement provided by TF-HNN is inversely correlated with the complexity of the task-specific module.
Table 4: Node classification accuracy (%) and training time for HNNs and TF-HNN on Trivago.
@@
Table 5: Hyperlink prediction AUC (%) and training time for HNNs and TF-HNN on Trivago.
@@
Table 6: The impact from the TF-MP-Module of TF-HNN in node classification. The best result on each dataset is highlighted in bold font.
@@
Table 7: The impact from the weighted $\mathbf{S}$ of TF-HNN in node classification. The best result on each dataset is highlighted in bold font.
@@
TF-HNN with Weighted $\mathbf{S}$ 84.90 ± 1.55 74.75 ± 1.62 68.87 ± 4.57 75.57 ± 1.81 76.02
TF-HNN without Weighted $\mathbf{S}$ 86.54 ± 1.32 74.82 ± 1.67 70.42 ± 2.74 76.29 ± 1.99 77.02
Scalability of state-of-the-art models. We compare the scalability of our TF-HNN against the top two methods from previous experiments, Deep-HGNN and ED-HNN. Deep-HGNN uses a direct feature aggregation function, while ED-HNN uses an indirect feature aggregation function. We evaluate the performance of all methods on node classification and hyperlink prediction using the large-scale hypergraph benchmark, Trivago, with the results detailed in Table 5 and Table 5, respectively. The results show that TF-HNN not only surpasses its competitors in performance but also dramatically reduces training time. Remarkably, in node classification, TF-HNN improves the accuracy by $10\%$ over the best baseline, Deep-HGNN, while using only $1\%$ of its training time. These findings underscore the scalability and value of TF-HNN for large-scale hypergraph applications. [^4] On Trivago, Deep-HGNN’s training time is much higher for node classification than hyperlink prediction. This is because a 64-layer model largely improves node classification accuracy, while a 2-layer model suffices for hyperlink prediction, as deeper models offer no benefit. The extra layers account for the longer training time.
@@ 6.3 Analysis
Impacts from the TF-MP-Module. In this study, we evaluate the effectiveness of our proposed TF-MP-Module within the TF-HNN framework by comparing it to two training-free message passing methods developed for graphs: SGC (Wu et al., 2019) and SIGN (Frasca et al., 2020). To adapt these methods to hypergraphs, we first project the given hypergraph into its clique expansion (CE), as introduced in Section 2, and then apply SGC and SIGN to this projected structure. Notably, we only replace the TF-MP-Module with SGC using CE and SIGN with CE, while keeping all other components of the TF-HNN unchanged. Table 6 show that our TF-MP-Module consistently outperforms the alternatives across all datasets. The superior performance of our method over SGC with CE is due to SGC with CE being a special case of our TF-MP-Module with $\alpha=0$, which, as discussed in Proposition 4.2, can cause node features to become overly similar to their neighbours, reducing distinctiveness. For SIGN with CE, the method concatenates features from different neighbor hops with the initial node features and processes them through a classifier, forcing it to handle both feature merging and classification, which increases complexity and can hinder convergence. In contrast, our TF-MP-Module directly merges neighbouring node information using the operator in Eq. (6), allowing the classifier to focus solely on classification, thereby simplifying the learning process and improving performance. These findings highlight the greater capacity of our TF-MP-Module compared to existing graph-based training-free modules in processing hypergraph structures, further underscoring its unique value for hypergraph machine learning.

### v6

@@ 3.3 Training-Free Message Passing
Learnable parameters removal. Based on the insights from Subsec. 3.2, to develop a training-free MP-Module, we first remove learnable parameters from the feature aggregation functions of the selected HNNs. Specifically, we replace the learnable matrices in these functions with identity matrices. This removal allows us to reformulate the feature aggregation functions as follows:
@@
Linearisation. Inspired by the insights from Subsec. 3.2, we further remove the non-linearity from the functions represented by Eqs. (4$\sim$4d), thereby transforming an MP-Module with $L$ feature aggregation functions into a more efficient single propagation step. For the sake of brevity, we present the simplification results as Proposition 3.2 and provide its mathematical proof in Appendix D.
@@
Let $\mathbf{X}_{\mathcal{V}}$ be input node features, $\hat{\mathbf{X}}_{\mathcal{V}}$ be the output of the MP-Module of an UniGCNII/AllDeepSets/ED-HNN/Deep-HGNN with $L$ MP layers, and $\alpha\!\in\![\!0,1\!)$. Assume that learnable parameters and the non-linearity are removed from the module. For each model, given a hypergraph =$\mathcal{H}\!=\!\!\{\!\mathcal{V},\!\mathcal{E},\!\mathbf{H}\!\}$, there exists a clique expansion $\mathcal{G}\!=\!\{\!\mathcal{V},\mathbf{W}\!\}$ to unify its output as the following formula $\hat{\mathbf{X}}_{\mathcal{V}}\!=\!\Big{(}(1\!-\!\alpha)^{L}\!\mathbf{W}^{L}\!+\!\alpha\!\sum_{l=0}^{L-1}(1\!-\!\alpha)^{l}\!\mathbf{W}^{l}\Big{)}\mathbf{X}_{\mathcal{V}}$.
This proposition shows that, via our simplification, the selected models can be unified into a single formula. Based on this observation, our TF-MP-Module is designed as the following equation:
@@
where $\delta(\cdot)$ is a function that returns 1 if $e_{k}$ connects $v_{i}$ and $v_{j}$ and returns 0 otherwise. We discuss the connection between the clique expansion mentioned in Proposition 3.2 and the one defined in Eq (5) in Appendix T. The main intuition behind the design of this edge weight is that, under certain conditions, the value of $\mathbf{W}_{H_{ij}}$ is positively correlated with the probability of nodes $v_{i}$ and $v_{j}$ having the same label. Consequently, this edge weight design can make the weighted clique expansion fit the homophily assumption, which is that connected nodes tend to be similar to each other (McPherson et al., 2001). Appendix F elaborates on this intuition. Furthermore, inspired by previous works in the graph domain (Wu et al., 2020), we generate $\mathbf{S}$ based on a symmetrically normalised $\mathbf{W}_{H}$ with the self-loop. Specifically, we generate the $\mathbf{S}$ by:
@@ 4 Theoretical Analysis
Let $H_{0}^{v^{L}_{i}}$ be the entropy of information [^3] We use “entropy of information” in a conceptual way - A detailed discussion is in Appendix U. used by an HNN with $L$ feature aggregation functions in generating the features of a node $v_{i}$, $H_{1}^{v^{L}_{i}}$ be the entropy of information used by an TF-HNN with an $L$-layer TF-MP-Module for the same purpose, and $H_{2}^{v^{L}_{i}}$ denote the entropy of information within node $v_{i}$ and its $L$-hop neighbours on the hypergraph. Then, $H_{0}^{v^{L}_{i}}=H_{1}^{v^{L}_{i}}=H_{2}^{v^{L}_{i}}$.
The proposition above theoretically demonstrates that TF-HNN is as effective as existing HNNs in utilising information from pre-defined hypergraphs to generate node features. Specifically, both TF-HNN and existing HNNs achieve this by aggregating neighbourhood information.
Robustness against the oversmoothing issue. According to Proposition 3.3, TF-HNN can use global interactions within the given hypergraph to generate node features by deepening the TF-MP-Module. However, as shown in a recent study (Chen et al., 2022), the message-passing-based models on hypergraphs may suffer from the oversmoothing issue. This issue refers to the tendency of a model to produce indistinguishable features for nodes in different classes as the model depth increases, which might degrade the performance of the framework in downstream tasks. Consequently, to further support the effectiveness of TF-HNN, we analyse its robustness against the oversmoothing issue. Inspired by works in the graph domain (Yang et al., 2021; Zhang et al., 2020; Singh et al., 2025), we show Proposition 4.2, with the proof provided in Appendix H.
@@
Based on Proposition 4.2, a deep TF-MP-Module tends to assign similar features to connected nodes on a hypergraph while maintaining the unique input information of each node. Hence, even with an extremely large number of layers, the TF-MP-Module can preserve the unique input information of each node, making it robust against the oversmoothing issue. This property further ensures the effectiveness of our TF-HNN. Sec. 6 empirically supports the analysis presented in this section.
@@ 5 Related Work
Recent efforts have leveraged hypergraphs to improve node feature learning for downstream tasks (Bick et al., 2023; Liu et al., 2024; Xu et al., 2024). The predominant models in this area are hypergraph neural networks (HNNs) with a message passing module that enables information exchange between connected nodes (Antelmi et al., 2023). However, as noted for their graph counterparts (Wu et al., 2019a; Frasca et al., 2020), HNNs suffer from low learning efficiency. Two recent works attempt to mitigate this limitation (Feng et al., 2024; Tang et al., 2024). In (Feng et al., 2024), an HNN named HGNN (Feng et al., 2019) is distilled into an MLP during training; while in (Tang et al., 2024), hypergraph structural information is integrated into an MLP via a loss function computed by the sum of the maximum node feature distance in each hyperedge. These approaches primarily focus on reducing model inference complexity via a dedicated training process. However, these designs still lead to hypergraph machine learning models with computationally intensive training procedures. For instance, the distillation process in (Feng et al., 2024) necessitates training a HNN with a computational complexity of $\mathcal{O}(M+TLm^{\prime}d+TLnd^{2})$, where with $M$ is the training computational complexity of the task-specific module, $n$ is the node count, $m^{\prime}$ is the number of edges in the clique expansion, $L$ is the number of layers, and $d$ is the feature dimension. Given that many existing models are applied to semi-supervised hypergraph node classification settings, which involve the simultaneous use of training and test data (Huang & Yang, 2021; Chien et al., 2022; Wang et al., 2023a; b; Joachims et al., 1999; Chapelle et al., 2009), reducing the training complexity of hypergraph-based models remains a critical yet underexplored challenge. We address this challenge with a novel model named TF-HNN, which is the first to decouple the processing of hypergraph structural information from model training.
@@ 6.1 Experimental Setup
Dataset and baseline. We conduct experiments on seven real-world hypergraphs: Cora-CA, DBLP-CA, Citeseer, Congress, House, Senate, which are from (Chien et al., 2022), and Trivago from (Kim et al., 2023). We use nine HNNs as baselines. Five of these baselines utilize direct feature aggregation, including HGNN (Feng et al., 2019), HCHA (Bai et al., 2021), UniGCNII (Huang & Yang, 2021), PhenomNN (Wang et al., 2023b), and Deep-HGNN (Chen et al., 2022). The remaining four baselines employ indirect feature aggregation, consisting of HNHN (Dong et al., 2020), AllDeepSets (Chien et al., 2022), AllSetTransformer (Chien et al., 2022), and ED-HNN (Wang et al., 2023a). The dataset statistics are summarized in Table 2, with detailed descriptions of datasets available in Appendix I.
@@
Implementation. For node classification, we follow previous works (Wang et al., 2023a; Duta et al., 2023) to use a 50%/25%/25% train/validation/test data split and adapt the baseline classification accuracy from them [^4] Since there are no reported results for Deep-HGNN under the chosen data split, we use our own results.. Additionally, similar to these works, we implement the classifier based on MLP for our TF-HNN and report the results from ten runs. For hyperlink prediction, existing works (Chen & Liu, 2023) primarily focus on the design of the prediction head, with no reported results for applying our baseline HNNs to this task, so we report all results from five runs conducted by ourselves. We employ the deep set function implemented by (Chien et al., 2022) as the prediction head for our method and baseline methods due to its simplicity. We use a 50%/25%/25% train/validation/test data split and ensure each split contains five times as many fake hyperedges as real hyperedges. TF-HNN and HNNs only use real hyperedges in training and validation sets for message passing. Experiments were on RTX 3090 GPUs by PyTorch. See our code here. More details are in Appendix L and  S.
@@ 6.2 Comparison with Baselines
Node classification. We summarize the classification accuracy and relative training time of the TF-HNN and HNNs in Table 3 and Figure 2, respectively. Table 3 demonstrates that TF-HNN not only leads to the best results across four datasets (Cora-CA, Congress, House, and Senate) but also results in the highest average mean accuracy overall. These findings confirm the effectiveness of TF-HNN in generating powerful node features for node classification. As discussed in Sec. 4, we attribute this effectiveness to the ability of TF-HNN to match the power of existing HNNs in utilising neighbourhood information to generate node features, while requiring the optimization of fewer parameters. This makes the TF-HNN more efficient in utilising training data and reduces the risk of overfitting, thereby outperforming more complex HNN counterparts. On the other hand, as shown in Figure 2, the TF-HNN consistently needs less training time to achieve superior performance compared to the HNNs. This highlights the high training efficiency of TF-HNN in downstream tasks.
@@
Hyperlink prediction. Figure 3 shows the results for hyperlink prediction, which exhibit a pattern similar to that observed in the hypergraph node classification task: the TF-HNN outperforms the HNNs while requiring less training time. An interesting observation is that the training efficiency improvement from TF-HNN is smaller in the hyperlink prediction task compared to the node classification task. The intuition is that the task-specific module used for hyperlink prediction is more computationally expensive than the one used for hypergraph node classification. Specifically, the hyperlink predictor needs to generate a prediction for each hyperedge by aggregating the features of the nodes connected by the hyperedge, whereas the node classifier only needs to perform forward propagation for each individual node. Appendix M discusses that the efficiency improvement provided by TF-HNN is inversely correlated with the complexity of the task-specific module.
Table 4: Node classification accuracy (%) and training time for models on Trivago.
@@
Table 5: Hyperlink prediction AUC (%) and training time for models on Trivago.
@@
Table 6: The impact from the TF-MP-Module in node classification. The best results are in bold font.
@@
Table 7: The impact from the weighted $\mathbf{S}$ in node classification. The best results are in bold font.
@@
TF-HNN without Weighted $\mathbf{S}$ 84.90 ± 1.55 74.75 ± 1.62 68.87 ± 4.57 75.57 ± 1.81 76.02
TF-HNN with Weighted $\mathbf{S}$ 86.54 ± 1.32 74.82 ± 1.67 70.42 ± 2.74 76.29 ± 1.99 77.02
Scalability of state-of-the-art models. We compare the scalability of our TF-HNN against the top two baseline methods from previous experiments, Deep-HGNN and ED-HNN. Deep-HGNN uses a direct feature aggregation function, while ED-HNN uses an indirect feature aggregation function. We evaluate the performance of all methods on node classification and hyperlink prediction using the large-scale hypergraph benchmark, Trivago, with the results detailed in Table 5 and Table 5, respectively. The results show that TF-HNN not only surpasses its competitors in performance but also dramatically reduces training time. Remarkably, in node classification, TF-HNN improves the accuracy by around $10\%$ over the best baseline, Deep-HGNN, while using only about $1\%$ of its training time. These findings underscore the value of TF-HNN for large-scale hypergraphs. [^5] On Trivago, Deep-HGNN’s training time is much higher for node classification than hyperlink prediction. This is because a 64-layer model largely improves node classification accuracy, while a 2-layer model suffices for hyperlink prediction, as deeper models offer no benefit. The extra layers account for the longer training time. Due to the space limit, additional discussions and experiments are in Appendix N, O, P, R, and V.
@@ 6.3 Analysis
Impacts from the TF-MP-Module. In this study, we evaluate the effectiveness of our proposed TF-MP-Module within the TF-HNN framework by comparing it to two training-free message passing methods developed for graphs: SGC (Wu et al., 2019a) and SIGN (Frasca et al., 2020). To adapt these methods to hypergraphs, we first project the given hypergraph into its clique expansion (CE), as introduced in Sec. 2, and then apply SGC and SIGN to this projected structure. Notably, we only replace the TF-MP-Module with SGC using CE and SIGN with CE, while keeping all other components of the TF-HNN unchanged. Table 6 show that our TF-MP-Module consistently outperforms the alternatives across all datasets. The superior performance of our method over SGC with CE is due to SGC with CE being a special case of our TF-MP-Module with $\alpha=0$, which, as discussed in Proposition 4.2, can cause node features to become overly similar to their neighbours, potentially reducing distinctiveness. For SIGN with CE, the method concatenates features from different neighbor hops with the initial node features and processes them through a classifier, forcing it to handle both feature merging and classification, which increases complexity and can hinder convergence. In contrast, our TF-MP-Module directly merges neighbouring node information using the operator in Eq. (6), allowing the classifier to focus solely on classification, thereby simplifying the learning process and improving performance. These findings highlight the greater capacity of our TF-MP-Module compared to existing graph-based training-free modules in processing hypergraph structures, further underscoring its unique value for hypergraph machine learning.